OpenStudy (harleyquinn98):

what is the soloution to the problem n over 18= 5 over 9. Find n A. 10 b 5 c. 2.5 d 15

OpenStudy (harleyquinn98):

@Shadowking625

OpenStudy (photon336):

\[\frac{ n }{ 18 } = \frac{ 5 }{ 9 }\]

OpenStudy (photon336):

we need to isolate n first by multiplying both sides by 18 like this. \[\frac{ n }{ 18 }*\frac{ 18 }{ 1 } = \frac{ 5 }{ 9 }*18\]

OpenStudy (harleyquinn98):

k

OpenStudy (photon336):

now as you can see, when we do this n is now by itself so now we're left with \[n = \frac{ 5 }{ 9 }*\frac{ 18 }{ 1 }\]

OpenStudy (anonymous):

Damn i just did that in my head lol

OpenStudy (harleyquinn98):

shut up im bad at math!!!!!!!

OpenStudy (anonymous):

Im amazing in math

OpenStudy (harleyquinn98):

ik thats why tagged u i txted u

OpenStudy (photon336):

so @HarleyQuinn98 all you need to do is multiply 18 by 5/9 \[\frac{ 5 }{ 9 }*\frac{ 18 }{ 1 } = \frac{ 90 }{ 9 } = 10\] and we can plug this back in to check our answer too n = 10 and we find that these two are equal. \[\frac{ 10 }{ 18 } = \frac{ 5 }{ 9 }\]

OpenStudy (anonymous):

Just cross multiply